Project

General

Profile

NDOS Commissioning Meeting 2011/03/10

Attendance: Jon, Enrique, Gavin, Mat, Mark Mathis, Chad, Andrew, Tricia, Alena, Susan, Leon, Ryan, Jarek, Minerba

Agenda:

  1. Installation/Channel Recovery Update
  2. DAQ Update & Plans
  3. DSO Studies - Update
  4. DCS Status & Plans
  5. Shifters' report
  6. AOB

Installation/Channel Recovery:

  • Mat:
    • all of the previously untested modules have been pressure tested. No leaks were found.
    • retested horizontals on block four
    • repair of cracked manifold in block 4 seems to have worked
  • Repair of secondary containment is next
  • Installation of muon catcher was completed earlier this week

DAQ Hardware:

  • Andrew:
    • work to increase frame size to 9k was completed earlier today around 2 pm.
    • an extra high performance network switch was found and will be installed on Monday.
  • Leon:
    • noise-cancellation board installation "tool" has not yet arrived at IU, hopefully tomorrow
    • Rick T. has lined up some technicians to do the board installation
  • A DCM (dcm-15) is now available for installation (from the test-stand)

DAQ Software:

  • Peter:
    • Testing of new version of DDS (underlying message system) system is underway; new version should reduce need to do frequent restarts of the message system
    • TDU issue: several DCMs at once (sometime all) suddenly start eating up all CPU. Possibly due to DCM getting out of whack in the timing chain, thus effecting other DCMs. Investigations are underway.
    • Ron Rechenmacher has done several studies to improve performance of DCMs (amount of data a DCM can push through without saturating the CPU usage). Latest change of MTU size from to 1.5 kB to 9 kB has ~doubled amount of data DCM can push through. Sue Kasahara is also looking into further improvements to DCMApp.
    • Occasionally we see "empty" events in the CR, because triggers seem to be lost between GlobalTrigger and BufferNodes. There are now some ideas to investigate the cause and remedy.

DSO Studies:

  • Mark Mathis:
    • moved plots to a new location (http://nusoft.fnal.gov/nova/dso)
    • a long DSO scan was done two days ago after the detector had been cooled. Analysis of these data is underway.

DCS:

  • Chad: RTDs arrived today and a module.
  • Athans: Sent display of sample output of his DCM DCS monitor program. FEB status register readback isn't quite working yet, but example code is in DCMProgramUtils. Question was asked about frequency of polling; Jon suggested that anywhere from 1-10 seconds should be fine.

Any other business: